OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (lynfran):

|2x+6|-4=20 first step add 4 to both sides of the equation we get |2x+6|-4+4=20+4 |2x+6|=20+4

OpenStudy (lynfran):

next step split the equation in 2...ie 1 positive and 1 negative ...like this |2x+6|=24 then (2x+6)=24 and (2x+6)=-24

OpenStudy (lynfran):

now we solve for x

OpenStudy (lynfran):

2x+6=24 2x=24-6 2x=18 x=18/2 x=9

OpenStudy (lynfran):

can u solve this one now 2x+6=-24

OpenStudy (anonymous):

2x+6=-24 = -9

OpenStudy (lynfran):

2x+6=-24 2x+6-6=-24-6 2x=-24-6 2x=-30 x=-30/2 x=-15

OpenStudy (anonymous):

so its -9 and -15 ???????

OpenStudy (anonymous):

how did you get -15

OpenStudy (lynfran):

2x+6=24 2x=24-6 2x=18 x=18/2 x=9 and 2x+6=-24 2x+6-6=-24-6 2x=-24-6 2x=-30 x=-30/2 x=-15

OpenStudy (anonymous):

ooh ok